Shift summary for Eve, 1/3/00


Subject: Shift summary for Eve, 1/3/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:30:52 CST


              Shift Summary for Eve Shift, 1/03/00

On Shift: A. Erwin, E. Blucher
           (+ Sasha Glazov, Rick K, Leo, Wah)
Hours of beam available at all intensities: 5
Hours of beam available at nominal intensity: 1
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...):3

Accounting of lost time:

========================================================================
Accesses during shift:

RC,EB - to NM2 to remove beam plug
EB,LB - to KTeV hall to move MA,Reg out of beam
RF,AE - to NM2 to remove shadow absorber.

========================================================================
Runs ended this shift:
15379 KAMI_LAMPI
15383 KAMI_LAMPI
15384 SHORT_REG832
15385 SHORT_REG832
15386 SHORT_REG832
15387 KAMI_LAMPI

========================================================================
Overview of shift:
  Special runs to prepare for KAMI test and 832 short reg. run

========================================================================
Description of Shift:

Start with shortened regenerator, 20" common absorber, shadow
absorber in beam, west beam hole plugged.

Run Spill Time
15379 15:30 First attempt at KAMI_LAMPI run. L3 code is crashing.
                    We spend most of the next couple of hours trying
                    to figure out what is causing code to crash.

                    We manage to take one good spill of data at 2E12.
                    Rick K estimates ~35 usable protons in hole per 2E12
                    (see KAMI logbook for details)

             17:10 Linac RF problem

             17:30 We take advantage of down time to make NM2 access to
                   remove beam plug.

             19:00 Beam returns

             19:15 Perform ke3 trigger study using reg + shadow
                   absorber as "beam plug". Trigger sheets with
                   ptkick = 400 MeV, 150 MeV in KAMI logbook I.

             19:15 Bug in Level 3 for KAMI_LAMPI found.

15383 19:25 Run in KAMI_LAMPI configuration (with 2 beams).
                   Level 3 is extremely slow because of very large
                   number of vertices.

             19:50 Request increase in intensity to 8e12.

15384 20:10 Start SHORT_REG832 run while intensity is increased.
 
15384 20:15 Run stopped with L3 accept problem. L1 trigger rates
                    for dibaryon trigger are very high (110 kHz for trig.
                    11)
15384 20:20 Beam intensity fluctuations are terrible, so we
                    close beam stop. There are aborts on every spill

             20:50 Waiting for beam. Add HA1 and HA2 to H-dibaryon triggers
                    to try to reduce rates.

             21:20 MCR still having no success tuning beam. They decide to
                    give up on two turns until they consult with experts.
                    We will continue special runs with 4e12.

15385 50 21:21 Beam stop opened.

             21:30 Lost quadrant of CsI. Stop run and do fresh start.

15386 21:40 Start SHORT_REG832 run. Still 20" common absorber,
                   shadow absorber. SEM=4e12.

             21:45 Close beam stop while MCR continues tuning. They believe
                   they've found the problem (a bad QXR parameter)

        13 21:57 Reopen beam stop

        15 22:05 Change common absorber to 10"; user latch = 2; sem=4e12

        22 Back to 20" common absorber; user latch=3; sem=8e12

15386 22:35 Stop run.

15387 22:45 One more KAMI_LAMPI run to test some new cuts.

             22:55 Turn off beam for NM2 cooldown prior to access to
                   remove shadow absorber.

             23:10 Turn off reg and MA HV. Access to remove MA and REG
                   from beam.

             23:30 Access to NM2 to remove shadow absorber.

15388 0:00 Start BEAM_799 run - waiting for beam.
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

The hot water heater for the building is not working.



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:30:55 CST